OpenStudy (anonymous):

Cuadratic equasion x+1/6= x-1/x

OpenStudy (anonymous):

multiply both sides of the equation by x to make the equation quadratic.

OpenStudy (anonymous):

\[x(x+\frac{1}{6})\ =\ x(x-\frac{1}{x})\]\[x^2+\frac{1}{6}x=x^2-1\]\[\frac{1}{6}x=-1\]\[x=-6\]